Circular No. 8153

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

VARIABLE STAR NEAR UGC 10700 Further to IAUC 8147, J. Graham and W. Li report the LOTOSS

discovery, on unfiltered KAIT images taken on June 17.4 and 18.4 UT, of a variable star (mag ∼ 16.5) located at α = 17^{h}06^{m}06^{s}.12, δ = +25^{o}51′53′′.3 (equinox 2000.0), which is 51′′.5 east and 79′′.9 south of the nucleus of UGC 10700. A KAIT image taken on June 11.4 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVA 2003ed IN NGC 5303A C. J. Stockdale, Naval Research Laboratory (NRL); R. A. Sramek,

National Radio Astronomy Observatory; S. D. Van Dyk, Infrared Processing and Analysis Center, California Institute of Technology; K. W. Weiler, NRL; and N. Panagia, European Space Agency and Space Telescope Science Institute, report the detection of radio emission from the type-II supernova 2003ed (IAUC 8129) with the Very Large Array. On June 17.12 UT in ‘A’ configuration, radio-flux densities of 1.81 ± 0.33 mJy at 22.46 GHz (wavelength 1.3 cm), 2.35 ± 0.26 mJy at 14.94 GHz (2.0 cm), 1.77 ± 0.11 mJy at 8.46 GHz (3.5 cm), and 0.65 ± 0.07 mJy at 4.86 GHz (6.2 cm) were measured at α = 13^{h}47^{m}45^{s}.39, δ = +38^{o}18′21′′.0 (equinox 2000.0; ± 0′′.2 in each coordinate). This is in close agreement with the reported optical position (end figures 45^{s}.36, 20′′.3; IAUC 8129). The radio emission is apparently still optically thick at 4.86 and 8.46 GHz and slightly more radio luminous in all bands, at a comparable age, than was SN 1993J (to which SN 2003ed has been compared; cf. IAUC 8144). Radio monitoring is continuing.

COMET P/2003 L1 (SCOTTI) Clearly diffuse NEAT images of this comet, taken with the Palomar

1.2-m Schmidt telescope on three nights in 2002 April, were identified and measured by M. Meyer. Additional astrometry and the following orbital elements (MPEC 2003-M21) confirm the suspicion (cf. IAUC 8145) that this is a short-period comet.

Epoch = 2003 Mar. 22.0 TT

T = 2003 Mar. 7.6754 TT  ω = 355.^{o}1927
e = 0.252002  Ω = 226.0386 }2000.0
q = 5.009723 AU  i = 9.0242

a = 6.697506 AU n^{o} = 0.0568636 P = 17.33 years

2003 June 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green